OpenStudy (lgbasallote):

Prove that if n is a perfect square, then n +2 is not a perfect square

OpenStudy (lgbasallote):

i suppose first step would be to assume n is perfect square

OpenStudy (lgbasallote):

so n = m^2

OpenStudy (lgbasallote):

then i suppose i find n^2

OpenStudy (lgbasallote):

i mean n + 2

OpenStudy (lgbasallote):

n + 2 would be m^2 + 2

OpenStudy (lgbasallote):

so now... i assume m^2 + 2 is a perfect square...

OpenStudy (anonymous):

we have no 2 perfect square numbers with difference 2

OpenStudy (lgbasallote):

so i let m^2 + 2 = k^2

OpenStudy (lgbasallote):

@mukushla that's not really a proof.....

OpenStudy (shubhamsrg):

you may continue @igba

OpenStudy (lgbasallote):

hmm i suppose the next step would be m^2 - k^2 = 2 so (m+k)(m-k) = 2 i don't think this proves anything...

OpenStudy (anonymous):

actually that is :) \[m^2-k^2=2\]\[(m-k)(m+k)=2\]and this is impossible

OpenStudy (lgbasallote):

why so?

OpenStudy (shubhamsrg):

well,,you've almost proved it.. now m and k both are integers/.

OpenStudy (shubhamsrg):

and its k^2 -m^2 = 2 and not otherwise..

OpenStudy (lgbasallote):

oh yes

OpenStudy (lgbasallote):

2 = (k + m)(k-m)

OpenStudy (lgbasallote):

now...integers huh...

OpenStudy (anonymous):

\[m+k>m-k\]so\[m+k=2\]so\[m=k=1\]but it gives\[m-k=0\]

OpenStudy (shubhamsrg):

you know k>m.. so from (k+m)(k-m)= 1*2 =>k+m =2 and k-m = 1 this doesnt have an integral solution..

OpenStudy (lgbasallote):

why do i know k > m again?

OpenStudy (shubhamsrg):

m^2 + 2 = k^2 both m and k are positive integers hence k>m

OpenStudy (shubhamsrg):

in layman language,, k^2 - m^2 >0

OpenStudy (lgbasallote):

when was it assumed that m and k are positive?

OpenStudy (shubhamsrg):

what else do you mean by perfect square ?

OpenStudy (lgbasallote):

i don't see how that relates.... the definition of a perfect square is x = k^2

OpenStudy (shubhamsrg):

4 is a perfect sq since its 2^2 i.e. square of a positive integer.. and likewise..

OpenStudy (anonymous):

@mukushla "we have no 2 perfect square numbers with difference 2" Because 0,1,4,9,6,5,6,9,4,1 etc....

OpenStudy (lgbasallote):

(-2)^2 is also 4...

OpenStudy (shubhamsrg):

you take the absolute value

OpenStudy (shubhamsrg):

i mean you have to..

OpenStudy (lgbasallote):

why?

OpenStudy (shubhamsrg):

see,, (k+m)(k-m) =2 either both are positive are both are negative,,since RHS is postive.. so even if you take -ve values,, negative * negative = positive ,,i.e. treated as absolute value..

OpenStudy (lgbasallote):

...i don't think this is a valid proof....

OpenStudy (shubhamsrg):

another explanation might me x is a perfect square because its y^2 x = y^2 = (-y)^2 sqrt(x) = | y | = | -y |

jhonyy9 (jhonyy9):

who can writing here now two perfect squares with difference of 2 ? so because i think do not exist or .... ???

OpenStudy (lgbasallote):

since when did sqrt x become |-y|

OpenStudy (anonymous):

0 to 9 squares end in 0,1,4,9,6,5,6,9,4,1 and pattern repeats. No +2 in there.

jhonyy9 (jhonyy9):

so and i think that this difference of two will be the ,,key" of this proof of ...

jhonyy9 (jhonyy9):

@lgbasallote what is your opinion from this ,please ?

OpenStudy (lgbasallote):

my opinion is that this is some tricky algebra

jhonyy9 (jhonyy9):

how do you think it ???

jhonyy9 (jhonyy9):

so and why ?

OpenStudy (lgbasallote):

because i can't think of a proof

OpenStudy (lgbasallote):

how can you prove/disprove that the difference of two squares is 2...

OpenStudy (lgbasallote):

hmmm...wait...

jhonyy9 (jhonyy9):

so i think this is very easy logicaly but to prove it will be difficile i think so but there are again @saifoo.khan and @satellite73 probably they will can doing it hope so much

OpenStudy (lgbasallote):

their fields of specialty aren't exactly in proving

OpenStudy (anonymous):

What's wrong with my proof?

OpenStudy (lgbasallote):

anyway... m^2 + 2 = k^2 k^2 - m^2 = 2 for it to be equal to 2...k^2 and m^2 should either be both even...or both odd so if i assume both are even.. then k and m are also both even so i can rewrite this as (2x)^2 - (2y^2) = 2 4x^2 - 4y^2 = 2 then... 2x^2 - 2y^2 = 1 x^2 - y^2 = 1/2 then since it's x^2 and y^2 are both integers...the difference should also be integer \('\therefore\) contradiction

OpenStudy (lgbasallote):

and i believe your proof @jhonyy9 did not involve proving...but brute force substitution..

OpenStudy (lgbasallote):

@estudier i mean

OpenStudy (lgbasallote):

@mukushla too

OpenStudy (anonymous):

So, that doesn't mean it's invalid..

OpenStudy (lgbasallote):

it does actually.....you can't do infinite substitutions to prove anything....

OpenStudy (anonymous):

There is no infinity involved, all integer squares end in the numbers I gave.

OpenStudy (lgbasallote):

infinite loops involves infinity agree?

OpenStudy (anonymous):

No infinite loop either, just logic.

OpenStudy (anonymous):

An integer ends in one of digits 0 to 9. The squares must end in etc....

OpenStudy (lgbasallote):

but we both know that setting values is illegal in proving

OpenStudy (anonymous):

I am not setting any values.

OpenStudy (lgbasallote):

and actually... you were checking 1^2, 2^2, 3^2, 4^2, etc. thus that will go on forever

OpenStudy (lgbasallote):

you're checking the squares of each number

OpenStudy (anonymous):

I am not.

OpenStudy (lgbasallote):

and though there is a pattern...there is still susbtitution involved

OpenStudy (lgbasallote):

and then you'll have to substitute forever to verify that the pattern really does not break

OpenStudy (lgbasallote):

then infinity is involved

OpenStudy (anonymous):

It's not a pattern, it's a fact 0 to 9 squares end in 0,1,4,9,6,5,6,9,4,1

OpenStudy (lgbasallote):

so you do admit that you checked the 0 to 9 squares

OpenStudy (anonymous):

Anything else is impossible.

OpenStudy (lgbasallote):

...that's substitution

OpenStudy (lgbasallote):

anyway...my point is...substitution isn't really allowed in proving

OpenStudy (anonymous):

You will find that this fact is used a lot in many valid proofs in number theory.

OpenStudy (lgbasallote):

perhaps

OpenStudy (anonymous):

OK, I can also do the algebra proof, but that doesn't mean it's any better than mine.

OpenStudy (lgbasallote):

since algebra is applicable in all numbers set to it....wouldn't that make it better than brute force?

OpenStudy (anonymous):

I am not using brute force, brute force is something like testing or experimental math. Or exhaustion (which is also a valid proof technique). It's the use of a fact about squares (ALL squares)

OpenStudy (lgbasallote):

can you prove it? that it's a fact about squares?

OpenStudy (anonymous):

Yes, all numbers end in 0 to 9 and so their squares end in .....QED.

OpenStudy (lgbasallote):

you didn't really demonstrate anything...you just repeated what you said

OpenStudy (anonymous):

You asked for proof, I just gave u a proof.

OpenStudy (anonymous):

If you deny my proof, provide a counterexample.

OpenStudy (lgbasallote):

since you try so hard to defend your proof, i'll accept it

OpenStudy (anonymous):

That's very gracious, thank you:-)

OpenStudy (lgbasallote):

do check on my algebra proof though please. i would like to know if that's the way contradiction works

OpenStudy (anonymous):

OK, I will look at it now....

OpenStudy (anonymous):

Did u do the case when both are odd? (Perhaps a more straightforward way is one you start with n= some k squared then next up must be (k+1)^2 = k^2 + 2k +1 and see what you can deduce from that...)

OpenStudy (lgbasallote):

why do i have to do odd too? if i already disproved using even...doesn't that prove by contradiction already?

OpenStudy (anonymous):

No because you divided the problem in cases, means you have to prove both (else you are missing all those cases)

OpenStudy (lgbasallote):

oh i see. but that kind of proof is right? because i'm a little skeptic if stating "2 can only be an outcome of even and even or odd and odd" is a valid statement

OpenStudy (lgbasallote):

again i ask... how is k > m

OpenStudy (anonymous):

(k+1)^2 = k^2 + 2k +1 >= n +3 > n+2 QED

OpenStudy (anonymous):

factoring involves lesser steps...its obvious that k>m \[k^2−m^2=2\]\[(k−m)(k+m)=2\]easily u can see \[k-m=1\]\[k+m=2\]add them\[2k=3\] Contradiction

OpenStudy (lgbasallote):

again i ask for the nth time @mukushla how is k > m

OpenStudy (anonymous):

ok man i just want to answer ur question

OpenStudy (anonymous):

"2 can only be an outcome of even and even or odd and odd" is a valid statement Why do you think it might not be valid?

OpenStudy (lgbasallote):

well because 8 - 4 is not 2 but it's even minus even